OpenStudy (sand-lock53):

i need help. very confusing. pls show me how to work it out. Luanne wants to have $3275 in her bank account in 10 years. How much money should she deposit if her account earns 7.5% interest which is compounded every month? A. $1589.01 B. $2159.45 C. $1921.53 D. $1550.62

OpenStudy (sooobored):

starting with basic definitions, there are 12 months in a year also, let x be equal to the amount of money we initially put in

OpenStudy (sooobored):

now, after every month, the bank will add an additional 7.5% in interest to the account any ideas how to represent this as an expression?

OpenStudy (sand-lock53):

is it (7.5% * 12) *10?

OpenStudy (sooobored):

or, what the bank account would look like after the first month?

OpenStudy (sooobored):

no

OpenStudy (sooobored):

what you have written makes no physical sense 10*12 would be the total amount of months but you cant apply interest to the total amount of months, you can only apply interest to a dollar amount

OpenStudy (sand-lock53):

oh ok

OpenStudy (sooobored):

ok, to rephrase, we start at month "0"- this is the initial amount of money we put into the account which would be x after 1 month, the bank adds an extra 7.5% to your bank account, so this means you still have x and you add x*(7.5%) so the total bank amount after 1 month is x+x*(7.5%)

OpenStudy (sooobored):

any questions thus far?

OpenStudy (sand-lock53):

no. what's next?

OpenStudy (sooobored):

quick quiz, if the initial amount of money i put in the bank is $100 how much money would i have after the first month?

OpenStudy (sand-lock53):

$10.75

OpenStudy (sooobored):

care to revise your answer? we're adding interest to $100 how can the resulting bank account be less than what we put in? did the bank steal some money?

OpenStudy (sand-lock53):

oh, sorry. I thoght the amount was $10. Anyways, it's $107.50

OpenStudy (sooobored):

ok, good, you understand how percentages work 7.5% is the equivalent of multiplying by 0.075 going back to teh expression we made, after 1 month, the total amount of money in the bank account would be x+0.075x

OpenStudy (sooobored):

adding up like terms, we get the expression 1.075 x after the first month

OpenStudy (sooobored):

any questions?

OpenStudy (sand-lock53):

I'm assuming x still equals zero?

OpenStudy (sooobored):

no, x is a variable we use to denote the total amount of money we initially put it

OpenStudy (sooobored):

it would be zero is we decided to put zero dollars into the bank account initially in our example, we initally put in 100 dollars it could be any of the multiple choice answers its just like a placeholder since we dont know the actual amount

OpenStudy (sooobored):

so we use x, in order to determine a hypothetical situation since we dont know enough information to start from the beginning

OpenStudy (sand-lock53):

got it. what now?

OpenStudy (sooobored):

ok, assuming after the first month, we have (1.075 x) dollars in our bank account after the 2nd month, we add an additional 7.5% to the bank account, any idea what this expression might look like?

OpenStudy (sooobored):

if this doesnt make sense, think back to our example after the first month, we have $107.5 in the account, and after 1 additional month, we add 7.5% interest of $107.5 to $107.50

OpenStudy (sand-lock53):

so, would it be (1.75x)?

OpenStudy (sooobored):

er, highly unlikely

OpenStudy (sooobored):

ok, ill walk you through this

OpenStudy (sooobored):

after month 1, we have 1.075x 7.5% of 1.075x would be 0.075*( 1.075x) we add this interest to the amount we have in the account so it looks like 1.075x +0.075(1.075x)

OpenStudy (sooobored):

or in our $100 example we start with $107.50 in our account after the first month 7.5% of 107.50 is $8.0625 adding the interest to our bank account 107.50 +8.0625 so we have $115.5625 after the 2nd month

OpenStudy (sooobored):

makes sense?

OpenStudy (sand-lock53):

so do i just keep adding 7.5% to the amount of money?

OpenStudy (sooobored):

that essentially what the problem means, but there's a trick

OpenStudy (sooobored):

before we get into the trick and before i answer " do i need to do this 120 times"

OpenStudy (sooobored):

we look back at the expression after 2 months 1.075x +0.075(1.075x) now, if you're at all familar with the distribution property, or adding "like terms" we can simplify this expression since 1*(1.075x) +0.075(1.075x) through distribution/factoring property (1+0.075)(1.075x) 1.075*(1.075x) now, we could multiply the two 1.075 together, but that wont help you understand the concept

OpenStudy (sooobored):

so that means at the end of the 2nd month, we will have 1.075*1.075x in our bank account?

OpenStudy (sooobored):

so one last month calculation before i tell you the answers to all your problems after the 3rd month, how much money would we have in our bank account, or what does the expression look like?

OpenStudy (sand-lock53):

$1.23

OpenStudy (sooobored):

er? assuming we started with $100 ? could you please show your steps since i would like to know at what step you are making a mistake

OpenStudy (sand-lock53):

oh, i thought you were on the original problem, my apologies. is the equation 1.075*1.075+(7.5%+11556.25x)?

OpenStudy (sand-lock53):

107.5*107.5 = 11556.25, and then 7.5%+11556.25x

OpenStudy (sooobored):

ok, 1st mistake its 1.075*1.075, not 107.5 *107.5

OpenStudy (sooobored):

and i told you to leave them seperate

OpenStudy (sooobored):

also, if you're going to multiply by a percentage, you usually use the decimal equivalent 1%= 0.01 10%=0.1 0.1% = 0.001 7.5% = 0.075

OpenStudy (sooobored):

when you want to find the percentage of something, you multiply the value by the decimal equivalent so, after the 3rd month, assuming we start with (1.075 *1.075 x) the interest would be 0.075*(1.075 *1.075 x) adding that interest onto the bank account (1.075 *1.075 x)+0.75*(1.075 *1.075 x)

OpenStudy (sooobored):

any questions?

OpenStudy (sand-lock53):

so far, none

OpenStudy (sooobored):

ok, using the method i describe for after determining month 2 can you simplify? (1.075 *1.075 x)+0.75*(1.075 *1.075 x)

OpenStudy (sand-lock53):

1.075*1.075x^2+0.75?

OpenStudy (sooobored):

er no, FACTORING / distribution soooo (1.075 *1.075 x)+0.075*(1.075 *1.075 x) lets assume that (1.075 *1.075 x)= u so we have u+0.075u 1 u + 0.075u (1+0.075) u 1.075 u so 1.075*(1.075 *1.075 x) when simplified

OpenStudy (sooobored):

ok so to compare them to each other initial month x 1st month 1.075x 2nd month 1.075*1.075 x 3rd month 1.075*1.075*1.075x if you see the pattern by now this means 4th would be 1.075*1.075*1.075*1.075 x ... 10th 1.075*1.075*1.075*1.075*1.075........... x

OpenStudy (sooobored):

and this is where exponents come in handy \[1.075 *1.075 = 1.075^2\] \[1.075*1.075*1.075=1.075^3\] \[1.075*1.075*1.075*1.075=1.075^4\] thus meaning at the 120th month our expression would look like \[1.075^{120}x\]

OpenStudy (sooobored):

does that make sense?

OpenStudy (sooobored):

ohhh, i think i made a biggg mistake

OpenStudy (sooobored):

instead of 7.5% every month, its 7.5/12 % every month its 7.5% interest yearly anyways simple amendment to the expression 7.5 / 12 =0 .625% hence we would use (1.00625)^120 x= 3275

OpenStudy (sooobored):

and then solve for x fyi 1.00625 ^ 120 = 2.112

OpenStudy (sand-lock53):

k. thx alot i appreciate it :)

OpenStudy (sooobored):

yup no problem